# MIH-TCCT: Mitigating Inconsistent Hallucinations in LLMs via Event-Driven Text-Code Cyclic Training

### v1

@@ 1 Introduction
Achieving human-like logical consistency in reasoning is essential for advancing artificial general intelligence (AGI) (Achiam et al., 2023; Pan et al., 2023). Despite significant advancements in large language models (LLMs) in addressing a variety of natural language processing tasks (Jiang et al., 2024; Mündler et al., ; Ribeiro et al., ), these models still tend to produce hallucinations due to their inability to maintain logical rigor (Zhang et al., 2023; Bao et al., 2024). These manifest in two main forms: 1) Input-conflicting hallucination, where LLMs generate content that diverges from the provided input (Dale et al., 2023; Zhong et al., 2021a); and 2) Context-conflicting hallucination, where LLMs create content that contradicts information they have previously generated (Pu et al., 2023; Liu et al., 2024), as shown in Figure 1. Given that these hallucinations may occur in various domains, there is an urgent need for a universal approach to effectively address these issues across different tasks.
@@
Recent studies have emphasized the effectiveness of synthetic datasets in mitigating in-consistent hallucinations related to specific tasks. For instance, synthetic mathematical datasets (Toshniwal et al., 2024; Huang et al., 2024) have shown reduced contradictions in the calculation process and significant improvements in solving mathematical problems. Additionally, some research has synthesized reasoning data in the logical reasoning (Nie et al., 2020; Saeed et al., 2021) field while providing the reasoning process (Dalvi et al., 2021) during synthesis. LOGIC-LM and SymbCOT (Pan et al., 2023; Xu et al., 2024) integrate LLMs with symbolic solvers to enhance logical problem-solving. Unfortunately, the synthetic datasets that have been developed are often tailored to specific tasks or domains, which can restrict their generalizability.
@@
Advancements in training LLMs on extensive code datasets, such as CoCoGen and CodeRL, have significantly reduced inconsistent hallucinations in code generation tasks (Madaan et al., 2022; Le et al., 2022). However, these improvements have not effectively transferred to general natural language tasks due to fundamental differences in semantics, language styles, and structures between code and natural language. The precision of code contrasts with the ambiguity of human language, posing challenges in adapting code-focused techniques for natural language understanding and generation. Recent approaches, such as program-of-thought (Chen et al., 2023) and program-assisted LMs (Gao et al., 2023), have sought to bridge this divide by interpreting natural language problems and generating corresponding code solutions. However, these methodologies fail to successfully transfer the logical consistency capabilities gained from code training to a broader spectrum of NLP tasks.
@@
Inspired by code-oriented LLMs that excel in generating logically consistent code, we propose that integrating code data is vital for enhancing the ability of LLMs to maintain logical consistency across various NLP applications. Furthermore, we observe that there exists a correspondence between the structure of event-based text and that of code. The event-based text can be transformed into structured code, utilizing programming constructs such as classes, objects, and functions, with the correspondence between the two structures shown in Figure 2. Thus, we introduce MIH-TCCT: Mitigating Inconsistent Hallucinations in LLMs via Event-Driven Text-Code Cyclic Training. This method facilitates the cyclic training of code and the original text, continuously aligning the stylistic characteristics of the two ”languages” to improve dual generation capabilities. Ultimately, this approach aims to transfer the logical rigor of code data into natural language, fundamentally enhancing the logical consistency of LLM outputs. This novel integration not only fosters more reliable reasoning processes, but also paves the way for developing AI systems capable of delivering coherent and contextually accurate responses, representing a significant advancement in the capabilities of language models.
@@ 2.1 Synthetic Data to Reduce Hallucinations
Recent investigations indicate that hallucination phenomena in large language models (LLMs) may stem from inherent issues within the training data. This highlights the necessity for high-quality reasoning datasets to mitigate these inconsistencies (Zhang et al., 2023; Jiang et al., 2024). Notable examples include LOGIQA (Liu et al., 2023), a dataset of logical reasoning derived from the Chinese Civil Service Examination; RECLOR (Yu et al., ) and AR-LSAT (Zhong et al., 2021b), based on standardized graduate admissions exams; and FOLIO (Han et al., 2022), a dataset featuring first-order logic annotations. However, the resource-intensive and time-consuming nature of their creation limits the scalability and accessibility of training applications.
As a result, researchers are increasingly exploring synthetic data as a viable alternative (Achiam et al., 2023; Dubey et al., 2024). Significant synthetic mathematical datasets include KPDDS, OpenMathInstruct-1, and MathGenie (Toshniwal et al., 2024; Huang et al., 2024; Lu et al., 2024). KPDDS synthesizes question-answer pairs using key points and exemplars, yielding KPMath, which encompasses over one million pairs. OpenMathInstruct-1 comprises 1.8 million problem-solution pairs synthesized from code-interpreter solutions for GSM8K and MATH benchmarks via the Mixtral model. MathGenie enhances a seed dataset by generating new questions through a back-translation approach. Despite these advancements, synthetic datasets are often tailored to specific tasks, particularly mathematics, which may restrict their generalizability and performance across diverse real-world applications. This underscores the necessity for further research to broaden their applicability.
@@ 2.2 COT Data to Reduce Hallucinations
Based on the investigation of synthetic datasets, Chain-of-Thought (CoT) strategies have also reduced hallucinations during reasoning tasks (Wei et al., 2022). Some research has synthesized reasoning data in the field of logical reasoning (Nie et al., 2020; Saeed et al., 2021) while providing the reasoning process (Dalvi et al., 2021) during synthesis. For example, LOGIC-LM integrates LLMs with symbolic solvers, transforming natural language problems into symbolic formulations to minimize inconsistencies (Pan et al., 2023). Similarly, SymbCOT enhances CoT prompting by incorporating symbolic expressions and logical rules (Xu et al., 2024).
@@ 2.3 Code Data to Reduce Hallucinations
Recent research has shown that training large language models (LLMs) on code datasets can reduce in-consistent hallucinations. Notable models such as CoCoGen and CodeRL have enhanced performance in structured reasoning and code generation by leveraging the strict syntax and semantics inherent in programming languages (Madaan et al., 2022; Le et al., 2022). However, the benefits observed in code-related tasks do not readily extend to general natural language processing (NLP) tasks due to the fundamental semantic differences between the two domains.
Innovative approaches like Program of Thought (PoT) utilize LLMs, particularly Codex, to express reasoning processes as programs while offloading computational tasks to external systems (Chen et al., ). Despite this novel framework, PoT may encounter challenges in contexts where translating natural language into code is not straightforward, which can introduce potential errors. Similarly, Program-Aided Language Models (PAL) enable LLMs to decompose natural language problems into executable steps, delegating execution to environments such as Python interpreters (Gao et al., 2023). While this structural separation simplifies the model’s role, it also restricts its ability to engage in comprehensive reasoning, making it susceptible to errors during decomposition. Consequently, the effective transfer of logical consistency gained from code training to broader NLP applications remains an open challenge, highlighting the need for further research.
@@ 3.3 Text-Code Cylic Training
To provide a more intuitive understanding, Figure 4 presents an example of converting event-driven text into code. In our case, we extract essential details about Gajendrakumar Mitra and Sumathanath Ghosh, the co-founders of ”Mitra & Ghosh Publishers,” established on March 9, 1934. Leveraging this information, we define two classes: ‘Person‘, which represents individual entities, and ‘PublishingCompany‘, which encapsulates the characteristics of the publishing firm. Instances of these classes are created to accurately represent both the founders and the company, resulting in precise and high-quality code.
@@ 3.4 Quality Assessment of Generated Code
During iterative training, the code2text approach utilizes the original text as groundtruth, whereas the text2code process lacks a definitive ”correct” code as groundtruth. This discrepancy hinders the implementation of iterative training. Therefore, a dynamic quality assessment of the code generated in each iteration is essential. This assessment can provide reliable feedback for text2code training, thus ensuring its effectiveness in subsequent model training.
@@ 4.1 Settings
Model We assess the performance of several prominent base models, including Qwen2.5-7B(Yang et al., 2024), Llama 3.1 Instruct-8B(Dubey et al., 2024), and Ministra-8B-Instruct(Jiang et al., 2023). These models were selected based on their state-of-the-art capabilities in natural language processing tasks, providing a diverse range of architectures for our analysis.
Dataset We selected the wiki-40b-en dataset (Guo et al., 2020) as source data for training in the consistency improvement of the model. This is a large-scale text dataset based on the English Wikipedia website. We validate our method on two tasks: text summarization using the CNN/Daily Mail dataset (Chen et al., 2016), which is a widely recognized benchmark, and question answering (QA) with HaluEval (Li et al., 2023), which assesses hallucination performance.
Baseline We selected the following baselines: 1) Base Model: Each of the selected models is evaluated in its original, instruct versions; 2) Naive Prompting: This approach involves adding straightforward, consistent instructions to guide the models’ responses; 3) Supervised Fine-Tuning (SFT): Each base model undergoes SFT specific to the datasets used for the tasks; 4) SymbCOT (Xu et al., 2024): A recent method that reformulates chain-of-thought (CoT) reasoning as symbolic CoT to enhance consistency. The instructions for the 1) base and 2) prompt methods are shown in the Appendix A.5.
Metric For the summarization task, we utilize AlignScore (Zha et al., 2023) and consistency metric in UniEval(Zhong et al., 2022) to assess the phenomenon of inconsistent hallucination. In addition, we evaluate the summarization performance using its other three key metrics: Coherence, Fluency, and Relevance. For the QA task using the HaluEval dataset(Li et al., 2023), which is specifically designed to assess inconsistent hallucination, the actual performance of the QA task has not been evaluated. Instead, we use AlignScore(Zha et al., 2023) and Anah-V2(Gu et al., 2024) as metrics to investigate inconsistent hallucination, without focusing on task performance.
@@ A.2 The Prompt for Predicting Code from Text
      def    init   (self, first  name, last  name, title=””):
@@
gajendrakumar  mitra = Person(first  name=”Gajendrakumar”, last  name=”Mitra”, title=”Mr”)
sumathanath  ghosh = Person(first  name=”Sumathanath”, last  name=”Ghosh”)
@@
mitra  ghosh  publishers = PublishingCompany( name=”Mitra & Ghosh”, Publishers founders=[gajendrakumar  mitra,sumathanath  ghosh], established  date=”1934-03-09”)
@@ A.3 The Prompt for Predicting Text from Code
      def    init   (self, first  name, last  name, title=””):
@@
gajendrakumar  mitra = Person(first  name=”Gajendrakumar”, last  name=”Mitra”, title=”Mr”)
sumathanath  ghosh = Person(first  name=”Sumathanath”, last  name=”Ghosh”)
@@
mitra  ghosh  publishers = PublishingCompany( name=”Mitra & Ghosh”, Publishers founders=[gajendrakumar  mitra,sumathanath  ghosh], established  date=”1934-03-09”)
@@ A.4 The Prompt for Predicting Reorganized Text
text=f”mitra  ghosh  publishers.founders[0].title
@@ Appendix B Supplement to the Experimental Section
Table 5: The performance of the model based on the Llama3.1-Instruct architecture across varying proportions of mixed homogeneous data using the CNN-Daily Mail dataset. The performance assessment incorporates three crucial metrics: Coherence, Fluency, and Relevance. The ”Summ-CM” score is calculated by multiplying these three metrics together.

### v3

@@ Abstract
MIH-TCCT: Mitigating Inconsistent Hallucinations in LLMs via Event-Driven Text-Code Cyclic Training
Xinxin You1111,†, Xien Liu1111,†, Qixin Sun2222,†, Huan Zhang3333, Kaiyin Zhou4444, Shaohui Liu4444, Guoping Hu3333, Shijin Wang3333, Si Liu2222, Ji Wu1111 1Tsinghua University, Beijing, China  2Beihang University, Beijing, China 3iFLYTEK Research, Beijing, China  4Beijing University of Posts and Telecommunications, Beijing, China yxx23@mails.tsinghua.edu.cn
@@ 1 Introduction
Achieving human-like logical consistency in reasoning is essential for advancing artificial general intelligence (AGI) Achiam et al. (2023); Pan et al. (2023). Despite significant advancements in large language models (LLMs) in addressing a variety of natural language processing tasks Jiang et al. (2024); Mündler et al. ; Ribeiro et al. , these models still tend to produce hallucinations due to their inability to maintain logical rigor Zhang et al. (2023); Bao et al. (2024). These manifest in two main forms: 1) Input-conflicting hallucination, where LLMs generate content that diverges from the provided input Dale et al. (2023); Zhong et al. (2021a); and 2) Context-conflicting hallucination, where LLMs create content that contradicts information they have previously generated Pu et al. (2023); Liu et al. (2024), as shown in Figure 1. Given that these hallucinations may occur in various domains, there is an urgent need for a universal approach to effectively address these issues across different tasks.
@@
Recent studies have emphasized the effectiveness of synthetic datasets in mitigating in-consistent hallucinations related to specific tasks. For instance, synthetic mathematical datasets Toshniwal et al. (2024); Huang et al. (2024) have shown reduced contradictions in the calculation process and significant improvements in solving mathematical problems. Additionally, some research has synthesized reasoning data in the logical reasoning Nie et al. (2020); Saeed et al. (2021) field while providing the reasoning process Dalvi et al. (2021) during synthesis. LOGIC-LM and SymbCOT Pan et al. (2023); Xu et al. (2024) integrate LLMs with symbolic solvers to enhance logical problem-solving. Unfortunately, the synthetic datasets that have been developed are often tailored to specific tasks or domains, which can restrict their generalizability.
@@
Advancements in training LLMs on extensive code datasets, such as CoCoGen and CodeRL, have significantly reduced inconsistent hallucinations in code generation tasks Madaan et al. (2022); Le et al. (2022). However, these improvements have not effectively transferred to general natural language tasks due to fundamental differences in semantics, language styles, and structures between code and natural language. The precision of code contrasts with the ambiguity of human language, posing challenges in adapting code-focused techniques for natural language understanding and generation. Recent approaches, such as program-of-thought Chen et al. (2023) and program-assisted LMs Gao et al. (2023), have sought to bridge this divide by interpreting natural language problems and generating corresponding code solutions. However, these methodologies fail to successfully transfer the logical consistency capabilities gained from code training to a broader spectrum of NLP tasks.
@@
Inspired by code-oriented LLMs that excel in generating logically consistent code, we propose that integrating code data is vital for enhancing the ability of LLMs to maintain logical consistency across various NLP applications. Furthermore, we observe that there exists a correspondence between the structure of event-based text and that of code. The event-based text can be transformed into structured code, utilizing programming constructs such as classes, objects, and functions, with the correspondence between the two structures shown in Figure 2. Thus, we introduce MIH-TCCT: Mitigating Inconsistent Hallucinations in LLMs via Event-Driven Text-Code Cyclic Training. This method facilitates the cyclic training of code and the original text, continuously aligning the stylistic characteristics of the two "languages" to improve dual generation capabilities. Ultimately, this approach aims to transfer the logical rigor of code data into natural language, fundamentally enhancing the logical consistency of LLM outputs. This novel integration not only fosters more reliable reasoning processes, but also paves the way for developing AI systems capable of delivering coherent and contextually accurate responses, representing a significant advancement in the capabilities of language models.
@@ 2.1 Synthetic Data to Reduce Hallucinations
Recent investigations indicate that hallucination phenomena in large language models (LLMs) may stem from inherent issues within the training data. This highlights the necessity for high-quality reasoning datasets to mitigate these inconsistencies Zhang et al. (2023); Jiang et al. (2024). Notable examples include LOGIQA Liu et al. (2023), a dataset of logical reasoning derived from the Chinese Civil Service Examination; RECLOR Yu et al. and AR-LSAT Zhong et al. (2021b), based on standardized graduate admissions exams; and FOLIO Han et al. (2022), a dataset featuring first-order logic annotations. However, the resource-intensive and time-consuming nature of their creation limits the scalability and accessibility of training applications.
As a result, researchers are increasingly exploring synthetic data as a viable alternative Achiam et al. (2023); Dubey et al. (2024). Significant synthetic mathematical datasets include KPDDS, OpenMathInstruct-1, and MathGenie Toshniwal et al. (2024); Huang et al. (2024); Lu et al. (2024). KPDDS synthesizes question-answer pairs using key points and exemplars, yielding KPMath, which encompasses over one million pairs. OpenMathInstruct-1 comprises 1.8 million problem-solution pairs synthesized from code-interpreter solutions for GSM8K and MATH benchmarks via the Mixtral model. MathGenie enhances a seed dataset by generating new questions through a back-translation approach. Despite these advancements, synthetic datasets are often tailored to specific tasks, particularly mathematics, which may restrict their generalizability and performance across diverse real-world applications. This underscores the necessity for further research to broaden their applicability.
@@ 2.2 COT Data to Reduce Hallucinations
Based on the investigation of synthetic datasets, Chain-of-Thought (CoT) strategies have also reduced hallucinations during reasoning tasks Wei et al. (2022). Some research has synthesized reasoning data in the field of logical reasoning Nie et al. (2020); Saeed et al. (2021) while providing the reasoning process Dalvi et al. (2021) during synthesis. For example, LOGIC-LM integrates LLMs with symbolic solvers, transforming natural language problems into symbolic formulations to minimize inconsistencies Pan et al. (2023). Similarly, SymbCOT enhances CoT prompting by incorporating symbolic expressions and logical rules Xu et al. (2024).
@@ 2.3 Code Data to Reduce Hallucinations
Recent research has shown that training large language models (LLMs) on code datasets can reduce in-consistent hallucinations. Notable models such as CoCoGen and CodeRL have enhanced performance in structured reasoning and code generation by leveraging the strict syntax and semantics inherent in programming languages Madaan et al. (2022); Le et al. (2022). However, the benefits observed in code-related tasks do not readily extend to general natural language processing (NLP) tasks due to the fundamental semantic differences between the two domains.
Innovative approaches like Program of Thought (PoT) utilize LLMs, particularly Codex, to express reasoning processes as programs while offloading computational tasks to external systems Chen et al. . Despite this novel framework, PoT may encounter challenges in contexts where translating natural language into code is not straightforward, which can introduce potential errors. Similarly, Program-Aided Language Models (PAL) enable LLMs to decompose natural language problems into executable steps, delegating execution to environments such as Python interpreters Gao et al. (2023). While this structural separation simplifies the model’s role, it also restricts its ability to engage in comprehensive reasoning, making it susceptible to errors during decomposition. Consequently, the effective transfer of logical consistency gained from code training to broader NLP applications remains an open challenge, highlighting the need for further research.
@@ 3.3 Text-Code Cylic Training
To provide a more intuitive understanding, Figure 4 presents an example of converting event-driven text into code. In our case, we extract essential details about Gajendrakumar Mitra and Sumathanath Ghosh, the co-founders of "Mitra & Ghosh Publishers," established on March 9, 1934. Leveraging this information, we define two classes: ‘Person‘, which represents individual entities, and ‘PublishingCompany‘, which encapsulates the characteristics of the publishing firm. Instances of these classes are created to accurately represent both the founders and the company, resulting in precise and high-quality code.
@@ 3.4 Quality Assessment of Generated Code
During iterative training, the code2text approach utilizes the original text as groundtruth, whereas the text2code process lacks a definitive "correct" code as groundtruth. This discrepancy hinders the implementation of iterative training. Therefore, a dynamic quality assessment of the code generated in each iteration is essential. This assessment can provide reliable feedback for text2code training, thus ensuring its effectiveness in subsequent model training.
@@ 4.1 Settings
Model We assess the performance of several prominent base models, including Qwen2.5-7BYang et al. (2024), Llama 3.1 Instruct-8BDubey et al. (2024), and Ministra-8B-InstructJiang et al. (2023). These models were selected based on their state-of-the-art capabilities in natural language processing tasks, providing a diverse range of architectures for our analysis.
Dataset We selected the wiki-40b-en dataset Guo et al. (2020) as source data for training in the consistency improvement of the model. This is a large-scale text dataset based on the English Wikipedia website. We validate our method on two tasks: text summarization using the CNN/Daily Mail dataset Chen et al. (2016), which is a widely recognized benchmark, and question answering (QA) with HaluEval Li et al. (2023), which assesses hallucination performance.
Baseline We selected the following baselines: 1) Base Model: Each of the selected models is evaluated in its original, instruct versions; 2) Naive Prompting: This approach involves adding straightforward, consistent instructions to guide the models’ responses; 3) Supervised Fine-Tuning (SFT): Each base model undergoes SFT specific to the datasets used for the tasks; 4) SymbCOT Xu et al. (2024): A recent method that reformulates chain-of-thought (CoT) reasoning as symbolic CoT to enhance consistency. The instructions for the 1) base and 2) prompt methods are shown in the Appendix A.5.
Metric For the summarization task, we utilize AlignScore Zha et al. (2023) and consistency metric in UniEvalZhong et al. (2022) to assess the phenomenon of inconsistent hallucination. In addition, we evaluate the summarization performance using its other three key metrics: Coherence, Fluency, and Relevance. For the QA task using the HaluEval datasetLi et al. (2023), which is specifically designed to assess inconsistent hallucination, the actual performance of the QA task has not been evaluated. Instead, we use AlignScoreZha et al. (2023) and Anah-V2Gu et al. (2024) as metrics to investigate inconsistent hallucination, without focusing on task performance.
@@ A.2 The Prompt for Predicting Code from Text
      def    init   (self, first  name, last  name, title=""):
@@
gajendrakumar  mitra = Person(first  name="Gajendrakumar", last  name="Mitra", title="Mr")
sumathanath  ghosh = Person(first  name="Sumathanath", last  name="Ghosh")
@@
mitra  ghosh  publishers = PublishingCompany( name="Mitra & Ghosh", Publishers founders=[gajendrakumar  mitra,sumathanath  ghosh], established  date="1934-03-09")
@@ A.3 The Prompt for Predicting Text from Code
      def    init   (self, first  name, last  name, title=""):
@@
gajendrakumar  mitra = Person(first  name="Gajendrakumar", last  name="Mitra", title="Mr")
sumathanath  ghosh = Person(first  name="Sumathanath", last  name="Ghosh")
@@
mitra  ghosh  publishers = PublishingCompany( name="Mitra & Ghosh", Publishers founders=[gajendrakumar  mitra,sumathanath  ghosh], established  date="1934-03-09")
@@ A.4 The Prompt for Predicting Reorganized Text
text=f"mitra  ghosh  publishers.founders[0].title
@@ Appendix B Supplement to the Experimental Section
Table 5: The performance of the model based on the Llama3.1-Instruct architecture across varying proportions of mixed homogeneous data using the CNN-Daily Mail dataset. The performance assessment incorporates three crucial metrics: Coherence, Fluency, and Relevance. The "Summ-CM" score is calculated by multiplying these three metrics together.